# Highest weight category structures on Rep(B) and full exceptional collections on generalized flag varieties over $\mathbb{Z}$

### v2

@@ 13.1 Generating ${\mathrm{rep}}(P)$
13.2 The set $\{\hat{M}_{v}\}_{v\in W^{P}}$
For $P$-dominant $\lambda$ we use ^$\hat{\nabla}_{\lambda}$ as another notation for ${\mathrm{ind}}_{B}^{P}{\sf k}_{\lambda}$. Let us be given a set $\{\hat{M}_{v}\}_{v\in W^{P}}$ of objects of ${\mathrm{rep}}_{\mathrm{fr}}(P)$ with the following properties. For each ${v\in W^{P}}$ there is a diagram of $P$-modules
@@ 13.2 The set $\{\hat{M}_{v}\}_{v\in W^{P}}$
Examples of possible choices of $\hat{M}_{v}$ are ^$\hat{\nabla}_{{e_{v}}}$ with $f=g={\mathrm{id}}$, ${\mathrm{ind}}_{B}^{P}Q(e_{v})$ with $g={\mathrm{id}}$, $P(-e_{v})^{*}$ with $f={\mathrm{id}}$. The multiplicity of the weight $e_{v}$ is one in $\hat{M}_{v}$, $\hat{N}_{v}$, ^$\hat{\nabla}_{e_{v}}$. Every weight $\lambda$ of $\hat{M}_{v}$ satisfies $\lambda\leq_{a}e_{v}$.
@@
The smallest strictly full additive subcategory of ${\mathrm{rep}}(P)$ that
@@
is the category ${\mathrm{rep}}(P)$ of finite dimensional $P$-modules.
@@
Let $\tau\in X^{P}_{+}$. Let $V\in{\mathrm{rep}}(P)$ such that every weight $\mu$ of $V$ satisfies $\mu<_{a}\tau$. Then $V$ is an object of the smallest strictly full additive subcategory ${\mathcal{S}}$ of ${\mathrm{rep}}(P)$ satisfying
  • ${\mathcal{S}}$ contains $\hat{M}_{v}$ for $v\in W^{P}$ with $e_{v}<_{a}\tau$,
@@
We argue by induction along the well-ordered partial order $\leq_{a}$. So we assume the Theorem when $\tau$ is replaced by a $\sigma\in X_{+}^{P}$ with $\sigma<_{a}\tau$. Let $\sigma\in X_{+}^{P}$ with $\sigma<_{a}\tau$. Thus $V\in{\mathcal{S}}$ when all weights $\xi$ of $V$ satisfy $\xi<_{a}\sigma$.
@@
Step 1. First we wish to show that ${\mathcal{S}}$ contains at least one $P$-module $N$ with $N_{\sigma}={\sf k}_{\sigma}$ and with weights $\xi$ that satisfy $\xi\leq_{a}\sigma$. If $\sigma$ is a Steinberg weight $e_{v}$, then $v\in W^{P}$ by Lemma 13.8 and we simply take $N=\hat{M}_{v}$. If $\sigma$ is not a Steinberg weight, then we know from the proof Theorem 8.16 that there is a fundamental weight $\omega_{j}$ and a weight $\mu$ with $(\mu,\mu)<(\sigma,\sigma)$, so that the $B$-module $L=\nabla_{\omega_{j}}\otimes e^{\mu}$ has $\sigma$ as a weight of multiplicity one, and so that all weights $\nu$ of $L$ satisfy $\nu\leq_{a}\sigma$. Now take $N={\mathrm{ind}}_{B}^{P}L=\nabla_{\omega_{j}}\otimes{\mathrm{ind}}_{B}^{P}e^{\mu}$. By Lemma 13.9 every weight $\nu$ of $N$ satisfies $\nu\leq_{a}\sigma$ and $N_{\sigma}={\sf k}_{\sigma}$. In particular, ${\mathrm{ind}}_{B}^{P}e^{\mu}$ is nonzero, so $\mu\in X^{P}_{+}$ and ${\mathrm{ind}}_{B}^{P}e^{\mu}=\hat{\nabla}_{\mu}$. And every weight $\nu$ of ^$\hat{\nabla}_{\mu}$ satisfies $(\nu,\nu)<(\sigma,\sigma)$, hence $\nu<_{a}\sigma$, so ^$\hat{\nabla}_{\mu}$ lies in ${\mathcal{S}}$ and $N=\nabla_{\omega_{j}}\otimes\hat{\nabla}_{\mu}$ lies in ${\mathcal{S}}$.
@@
Step 2. Next we wish to show that ^$\hat{\nabla}_{\sigma}$ lies in ${\mathcal{S}}$. To this end we look for more $P$-modules $N$ in ${\mathcal{S}}$ with $N_{\sigma}={\sf k}_{\sigma}$ and such that the weights $\xi$ of $N$ satisfy $\xi\leq_{a}\sigma$. Start with the $N$ from Step 1. Let $N^{1}$ be the span of the weight spaces $N_{\mu}$ with $(\mu,\omega_{\alpha})<(\sigma,\omega_{\alpha})$ for at least one $\alpha\in\Pi^{\notin P}$. Then $N^{1}$ is a $P$-submodule that lies in ${\mathcal{S}}$. So we may replace $N$ with $N/N^{1}$ and further assume $N^{1}=0$. Let $N^{2}$ be the span of the weight spaces $N_{\mu}$ with $(\mu,\omega_{\alpha})\leq(\sigma,\omega_{\alpha})$ for all $\alpha\in\Pi^{\notin P}$. Then $N^{2}$ is a $P$-submodule and $N/N^{2}$ lies in ${\mathcal{S}}$. So we may replace $N$ with $N^{2}$ and further assume $N^{1}=0$, $N=N^{2}$. Let $N^{3}$ be the $P$-submodule of $N$ generated by $N_{\sigma}$. Then $N/N^{3}$ is in ${\mathcal{S}}$, so we may replace $N$ with $N^{3}$. Now the unipotent radical of $P$ acts trivially on $N$ and the projection $p$ of $N$ onto its weight space $N_{\sigma}$ is $T$-equivariant and $G_{P}$-equivariant, hence $P$-equivariant. By Lemma 13.9 the map $N\to\hat{\nabla}_{\sigma}$ induced by $p$ has kernel and cokernel in ${\mathcal{S}}$. So $\nabla_{\sigma}$ lies in ${\mathcal{S}}$.
@@
Step 3. Finally we want to show that ${\mathcal{S}}$ contains every $P$-module $V$ all whose weights $\mu$ satisfy $\mu<_{a}\tau$. Consider such a $V$. We may and shall assume that ${\mathcal{S}}$ contains every $P$-module whose set of weights is a proper subset of the set of weights of $V$. Say $V$ is nonzero. Choose an extremal weight $\sigma$ of $V$ that is $P$-dominant. Let $V^{1}$ be the span of the weight spaces $V_{\mu}$ with $(\mu,\omega_{\alpha})<(\sigma,\omega_{\alpha})$ for at least one $\alpha\in\Pi^{\notin P}$. This is a $P$-submodule, and if $V^{1}$ is nonzero, then $V$ is in ${\mathcal{S}}$ because both $V^{1}$ and $V/V^{1}$ are. Let $V^{2}$ be the span of the weight spaces $V_{\mu}$ with $(\mu,\omega_{\alpha})\leq(\sigma,\omega_{\alpha})$ for all $\alpha\in\Pi^{\notin P}$. This is a $P$-submodule, and if $V^{2}\not=V$, then $V$ is in ${\mathcal{S}}$ because both $V^{2}$ and $V/V^{2}$ are. So we further assume $V^{1}=0$ and $V=V^{2}$. Then the unipotent radical of $P$ acts trivially on $V$. Let $V^{3}$ be the $P$-submodule generated by $V_{\sigma}$. If $V^{3}\not=V$, then $V$ is in ${\mathcal{S}}$ because both $V^{3}$ and $V/V^{3}$ are. So we may assume $V=V^{3}$. The weights $\xi$ of $V$ now satisfy $\xi\leq_{a}\sigma$. The projection $p$ of $V$ onto its weight space $V_{\sigma}$ is $T$-equivariant and $G_{P}$-equivariant, hence $P$-equivariant. We have ${\mathrm{ind}}_{B}^{P}(V_{\sigma})=\hat{\nabla}_{\sigma}\otimes(V_{\sigma})_{\mathrm{triv}}$, where $(V_{\sigma})_{\mathrm{triv}}$ is $V_{\sigma}$ with trivial $P$-action, cf. [23, Proposition 17]. Consider the map $f:V\to{\mathrm{ind}}_{B}^{P}(V_{\sigma})=\hat{\nabla}_{\sigma}\otimes(V_{\sigma})_{\mathrm{triv}}$, corresponding with $p$. Both ${\mathsf{ker}}(f)$ and ${\mathsf{coker}}(f)$ lie in ${\mathcal{S}}$. As ^$\hat{\nabla}_{\sigma}$ lies in ${\mathcal{S}}$, so does ^$\hat{\nabla}_{\sigma}\otimes(V_{\sigma})_{\mathrm{triv}}$. So $V$ lies in ${\mathcal{S}}$. ∎
For completeness we mention
Corollary 13.12.
Let $\lambda\in X_{+}^{P}$. Then [^$[\hat{\nabla}_{\lambda}]$ lies in the $R(G)$-submodule of $R(P)$ generated by the $[\hat{M}_{v}]$ with $v\in W^{P},\;e_{v}\leq_{a}\lambda$.
Proof.
By induction along $\leq_{a}$. Either $\lambda$ is a Steinberg weight or it is not …∎
Definition 24.
For $v\in W^{P}$ put $\hat{Q}(e_{v})={\mathrm{ind}}_{B}^{P}Q(e_{v})$. And then for $p\in W^{P}$
  • $\hat{\sf Q}_{\succeq p}:={\mathrm{hull}}(\{\nabla_{\lambda}\otimes\hat{Q}(e_{v})\}_{v\succeq p,v\in W^{P},\lambda\in X(T)_{+}})$
  • $\hat{\sf Q}_{\succ p}:={\mathrm{hull}}(\{\nabla_{\lambda}\otimes\hat{Q}(e_{v})\}_{v\succ p,v\in W^{P},\lambda\in X(T)_{+}})$
  • $\hat{\sf P}_{\preceq p}:={\mathrm{hull}}(\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\preceq p,v\in W^{P},\lambda\in X(T)_{+}})$
  • $\hat{\sf P}_{\prec p}:={\mathrm{hull}}(\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\prec p,v\in W^{P},\lambda\in X(T)_{+}})$
@@
Given a $p\in W^{P}$, the triangulated hull in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits({\mathrm{Rep}}(P))$ of the two categories $\hat{\sf P}_{\preceq p}$ and $\hat{\sf Q}_{\succ p}$ is $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))$.
@@
    $\mathrm{R}^{i}{\mathrm{ind}}_{P}^{G}(P(-e_{v})\otimes\hat{Q}(e_{v}))=\begin{cases}{\sf k}&\text{if $i=0$,}\\ 0&\text{else.}\end{cases}$
@@
    If $w\not\leq v$ then $\mathrm{R}^{i}{\mathrm{ind}}_{P}^{G}(P(-e_{v})\otimes\hat{Q}(e_{w}))=0$ for all $i$.
@@
By [46, Theorem 1.9.(a)(ii)] the module $Q(e_{w})$ is ${\mathrm{ind}}_{B}^{P}$ acyclic. We get
$\mathrm{R}^{j}{\mathrm{ind}}_{B}^{P}(P(-e_{v})\otimes Q(e_{w}))=P(-e_{v})\otimes\mathrm{R}^{j}{\mathrm{ind}}_{B}^{P}(Q(e_{w}))=\begin{cases}P(-e_{v})\otimes\hat{Q}(e_{w})&\text{if $j=0$,}\\ 0&\text{else.}\end{cases}$
So $\mathrm{R}^{i}{\mathrm{ind}}_{B}^{G}(P(-e_{v})\otimes Q(e_{w}))=\mathrm{R}^{i}{\mathrm{ind}}_{P}^{G}(P(-e_{v})\otimes\hat{Q}(e_{w}))$. But the left hand side is known from Theorem 6.6. ∎
With these replacements we now proceed as in the case $P=B$: For $p\in W^{P}$, let $\hat{X}_{p}$ be the image of $P(-e_{p})^{*}$ under the left adjoint of the inclusion of $\hat{\sf Q}_{\succeq p}$ into $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))$. And let $\hat{Y}_{p}$ be the image of $\hat{Q}(e_{p})$ under the right adjoint of the inclusion of $\hat{\sf P}_{\preceq p}$ into $\mathcal{D}$.
Let $f_{p}$ be the natural map from $P(-e_{p})^{*}$ to $Q(e_{p})$. Recall that $f_{p}$ factors as a surjection $P(-e_{p})^{*}\to{\sf k}_{e_{p}}$, followed by an injection ${\sf k}_{e_{p}}\to Q(e_{p})$. Let $\hat{f}_{p}:P(-e_{p})^{*}\to\hat{Q}(e_{p})$ be induced by $f_{p}$. Then $\hat{f}_{p}$ induces an isomorphism of $T$-modules $(P(-e_{p})^{*})_{e_{p}}\to(\hat{Q}(e_{p}))_{e_{p}}$ and the weights $\mu$ of ${\mathsf{ker}}(\hat{f}_{p})$, ${\mathsf{coker}}(\hat{f}_{p})$ satisfy $\mu<_{a}e_{p}$. We have an exact triangle

### v3

@@ 13.1 Generating ${\mathrm{rep}}(\mathbf{P})$
13.2 The set $\{\hat{M}_{v}\}_{v\in W^{\mathbf{P}}}$
For $\mathbf{P}$-dominant $\lambda$ we use ^$\hat{\nabla}_{\lambda}$ as another notation for ${\mathrm{ind}}_{\mathbf{B}}^{\mathbf{P}}{\sf k}_{\lambda}$. Let us be given a set $\{\hat{M}_{v}\}_{v\in W^{\mathbf{P}}}$ of objects of ${\mathrm{rep}}_{\mathrm{fr}}(\mathbf{P})$ with the following properties. For each ${v\in W^{\mathbf{P}}}$ there is a diagram of $\mathbf{P}$-modules
@@ 13.2 The set $\{\hat{M}_{v}\}_{v\in W^{\mathbf{P}}}$
Examples of possible choices of $\hat{M}_{v}$ are ^$\hat{\nabla}_{{e_{v}}}$ with $f=g={\mathrm{id}}$, ${\mathrm{ind}}_{\mathbf{B}}^{\mathbf{P}}Q(e_{v})$ with $g={\mathrm{id}}$, $P(-e_{v})^{*}$ with $f={\mathrm{id}}$.
The multiplicity of the weight $e_{v}$ is one in $\hat{M}_{v}$, $\hat{N}_{v}$, ^$\hat{\nabla}_{e_{v}}$. Every weight $\lambda$ of $\hat{M}_{v}$ satisfies $\lambda\leq_{a}e_{v}$.
The following Theorem is similar to [3, Theorem 2], which is proved directly in the context of $\mathbf{G}$ and $\mathbf{P}$. Our argument differs in that it refers back to the proof for the $\mathbf{P}=\mathbf{B}$ case.
@@
The smallest strictly full additive subcategory of ${\mathrm{rep}}(\mathbf{P})$ that
@@
is the category ${\mathrm{rep}}(\mathbf{P})$ of finite dimensional $\mathbf{P}$-modules.
@@
Let $\tau\in X^{\mathbf{P}}_{+}$. Let $V\in{\mathrm{rep}}(\mathbf{P})$ such that every weight $\mu$ of $V$ satisfies $\mu<_{a}\tau$. Then $V$ is an object of the smallest strictly full additive subcategory ${\mathcal{S}}$ of ${\mathrm{rep}}(\mathbf{P})$ satisfying
  • ${\mathcal{S}}$ contains $\hat{M}_{v}$ for $v\in W^{\mathbf{P}}$ with $e_{v}<_{a}\tau$,
@@
We argue by induction along the well-ordered partial order $\leq_{a}$. So we assume the Theorem when $\tau$ is replaced by a $\sigma\in X_{+}^{\mathbf{P}}$ with $\sigma<_{a}\tau$. Let $\sigma\in X_{+}^{\mathbf{P}}$ with $\sigma<_{a}\tau$. Thus $V\in{\mathcal{S}}$ when all weights $\xi$ of $V$ satisfy $\xi<_{a}\sigma$.
@@
Step 1. First we wish to show that ${\mathcal{S}}$ contains at least one $\mathbf{P}$-module $N$ with $N_{\sigma}={\sf k}_{\sigma}$ and with weights $\xi$ that satisfy $\xi\leq_{a}\sigma$. If $\sigma$ is a Steinberg weight $e_{v}$, then $v\in W^{\mathbf{P}}$ by Lemma 13.9 and we simply take $N=\hat{M}_{v}$. If $\sigma$ is not a Steinberg weight, then by Corollary 8.24 there is a fundamental weight $\omega_{j}$ and a weight $\mu$ with $(\mu,\mu)<(\sigma,\sigma)$, so that the $\mathbf{B}$-module $L=\nabla_{\omega_{j}}\otimes e^{\mu}$ has $\sigma$ as a weight of multiplicity one, and so that all weights $\nu$ of $L$ satisfy $\nu\leq_{a}\sigma$. Now take $N={\mathrm{ind}}_{\mathbf{B}}^{\mathbf{P}}L=\nabla_{\omega_{j}}\otimes{\mathrm{ind}}_{\mathbf{B}}^{\mathbf{P}}e^{\mu}$. By Lemma 13.10 every weight $\nu$ of $N$ satisfies $\nu\leq_{a}\sigma$ and $N_{\sigma}={\sf k}_{\sigma}$. In particular, ${\mathrm{ind}}_{\mathbf{B}}^{\mathbf{P}}e^{\mu}$ is nonzero, so $\mu\in X^{\mathbf{P}}_{+}$ and ${\mathrm{ind}}_{\mathbf{B}}^{\mathbf{P}}e^{\mu}=\hat{\nabla}_{\mu}$. And every weight $\nu$ of ^$\hat{\nabla}_{\mu}$ satisfies $(\nu,\nu)<(\sigma,\sigma)$, hence $\nu<_{a}\sigma$, so ^$\hat{\nabla}_{\mu}$ lies in ${\mathcal{S}}$ and $N=\nabla_{\omega_{j}}\otimes\hat{\nabla}_{\mu}$ lies in ${\mathcal{S}}$.
@@
Step 2. Next we wish to show that ^$\hat{\nabla}_{\sigma}$ lies in ${\mathcal{S}}$. To this end we look for more $\mathbf{P}$-modules $N$ in ${\mathcal{S}}$ with $N_{\sigma}={\sf k}_{\sigma}$ and such that the weights $\xi$ of $N$ satisfy $\xi\leq_{a}\sigma$. Start with the $N$ from Step 1. Let $N^{1}$ be the span of the weight spaces $N_{\mu}$ with $(\mu,\omega_{\alpha})<(\sigma,\omega_{\alpha})$ for at least one $\alpha\in\Pi^{\notin\mathbf{P}}$. Then $N^{1}$ is a $\mathbf{P}$-submodule that lies in ${\mathcal{S}}$. So we may replace $N$ with $N/N^{1}$ and further assume $N^{1}=0$. Let $N^{2}$ be the span of the weight spaces $N_{\mu}$ with $(\mu,\omega_{\alpha})\leq(\sigma,\omega_{\alpha})$ for all $\alpha\in\Pi^{\notin\mathbf{P}}$. Then $N^{2}$ is a $\mathbf{P}$-submodule and $N/N^{2}$ lies in ${\mathcal{S}}$. So we may replace $N$ with $N^{2}$ and further assume $N^{1}=0$, $N=N^{2}$. Let $N^{3}$ be the $\mathbf{P}$-submodule of $N$ generated by $N_{\sigma}$. Then $N/N^{3}$ is in ${\mathcal{S}}$, so we may replace $N$ with $N^{3}$. Now the unipotent radical of $\mathbf{P}$ acts trivially on $N$ and the projection $p$ of $N$ onto its weight space $N_{\sigma}$ is $\mathbf{T}$-equivariant and $\mathbf{G}_{\mathbf{P}}$-equivariant, hence $\mathbf{P}$-equivariant. By Lemma 13.10 the map $N\to\hat{\nabla}_{\sigma}$ induced by $p$ has kernel and cokernel in ${\mathcal{S}}$. So ^$\hat{\nabla}_{\sigma}$ lies in ${\mathcal{S}}$.
@@
Step 3. Finally we want to show that ${\mathcal{S}}$ contains every $\mathbf{P}$-module $V$ all whose weights $\mu$ satisfy $\mu<_{a}\tau$. Consider such a $V$. We may and shall assume that ${\mathcal{S}}$ contains every $\mathbf{P}$-module whose set of weights is a proper subset of the set of weights of $V$. Say $V$ is nonzero. Choose an extremal weight $\sigma$ of $V$ that is $\mathbf{P}$-dominant. Let $V^{1}$ be the span of the weight spaces $V_{\mu}$ with $(\mu,\omega_{\alpha})<(\sigma,\omega_{\alpha})$ for at least one $\alpha\in\Pi^{\notin\mathbf{P}}$. This is a $\mathbf{P}$-submodule, and if $V^{1}$ is nonzero, then $V$ is in ${\mathcal{S}}$ because both $V^{1}$ and $V/V^{1}$ are. Let $V^{2}$ be the span of the weight spaces $V_{\mu}$ with $(\mu,\omega_{\alpha})\leq(\sigma,\omega_{\alpha})$ for all $\alpha\in\Pi^{\notin\mathbf{P}}$. This is a $\mathbf{P}$-submodule, and if $V^{2}\not=V$, then $V$ is in ${\mathcal{S}}$ because both $V^{2}$ and $V/V^{2}$ are. So we further assume $V^{1}=0$ and $V=V^{2}$. Then the unipotent radical of $\mathbf{P}$ acts trivially on $V$. Let $V^{3}$ be the $\mathbf{P}$-submodule generated by $V_{\sigma}$. If $V^{3}\not=V$, then $V$ is in ${\mathcal{S}}$ because both $V^{3}$ and $V/V^{3}$ are. So we may assume $V=V^{3}$. The weights $\xi$ of $V$ now satisfy $\xi\leq_{a}\sigma$. The projection $p$ of $V$ onto its weight space $V_{\sigma}$ is $\mathbf{T}$-equivariant and $\mathbf{G}_{\mathbf{P}}$-equivariant, hence $\mathbf{P}$-equivariant. We have ${\mathrm{ind}}_{\mathbf{B}}^{\mathbf{P}}(V_{\sigma})=\hat{\nabla}_{\sigma}\otimes(V_{\sigma})_{\mathrm{triv}}$, where $(V_{\sigma})_{\mathrm{triv}}$ is $V_{\sigma}$ with trivial $\mathbf{P}$-action, cf. [26, Proposition 17]. Consider the map $f:V\to{\mathrm{ind}}_{\mathbf{B}}^{\mathbf{P}}(V_{\sigma})=\hat{\nabla}_{\sigma}\otimes(V_{\sigma})_{\mathrm{triv}}$, corresponding with $p$. Both ${\mathsf{ker}}(f)$ and ${\mathsf{coker}}(f)$ lie in ${\mathcal{S}}$. As ^$\hat{\nabla}_{\sigma}$ lies in ${\mathcal{S}}$, so does ^$\hat{\nabla}_{\sigma}\otimes(V_{\sigma})_{\mathrm{triv}}$. So $V$ lies in ${\mathcal{S}}$. ∎
Remark 13.13.
In Theorem 13.11 and Theorem 13.12 we deal with ${\mathrm{rep}}(\mathbf{P})$, even over ${\mathbb{Z}}$, but the Corollary is about $R(\mathbf{P})$ as in the paper [3]. Of course it is easier to work in $R(\mathbf{P})$, but basically the arguments are the same and we get the same understanding of Steinberg weights and a similar constructive decomposition of elements of $R(\mathbf{P})$ over $R(\mathbf{G})$ as in the proofs of Ananyevskiy.
Definition 13.14.
For $v\in W^{\mathbf{P}}$ put $\hat{Q}(e_{v})={\mathrm{ind}}_{\mathbf{B}}^{\mathbf{P}}Q(e_{v})$. And then for $p\in W^{\mathbf{P}}$
  • $\hat{\sf Q}_{\succeq p}:={\mathrm{hull}}(\{\nabla_{\lambda}\otimes\hat{Q}(e_{v})\}_{v\succeq p,v\in W^{\mathbf{P}},\lambda\in X(\mathbf{T})_{+}})$
  • $\hat{\sf Q}_{\succ p}:={\mathrm{hull}}(\{\nabla_{\lambda}\otimes\hat{Q}(e_{v})\}_{v\succ p,v\in W^{\mathbf{P}},\lambda\in X(\mathbf{T})_{+}})$
  • $\hat{\sf P}_{\preceq p}:={\mathrm{hull}}(\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\preceq p,v\in W^{\mathbf{P}},\lambda\in X(\mathbf{T})_{+}})$
  • $\hat{\sf P}_{\prec p}:={\mathrm{hull}}(\{\nabla_{\lambda}\otimes P(-e_{v})^{*}\}_{v\prec p,v\in W^{\mathbf{P}},\lambda\in X(\mathbf{T})_{+}})$
@@
Given a $p\in W^{\mathbf{P}}$, the triangulated hull in $\mathop{\mathrm{D}\kern 0.0pt}\nolimits({\mathrm{Rep}}(\mathbf{P}))$ of the two categories $\hat{\sf P}_{\preceq p}$ and $\hat{\sf Q}_{\succ p}$ is $\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(P))$.
@@
    $\mathrm{R}^{i}{\mathrm{ind}}_{\mathbf{P}}^{\mathbf{G}}(P(-e_{v})\otimes\hat{Q}(e_{v}))=\begin{cases}{\sf k}&\text{if $i=0$,}\\ 0&\text{else.}\end{cases}$
@@
    If $w\not\leq v$ then $\mathrm{R}^{i}{\mathrm{ind}}_{\mathbf{P}}^{\mathbf{G}}(P(-e_{v})\otimes\hat{Q}(e_{w}))=0$ for all $i$.
@@
By Theorem 1.8, it suffices to treat the case where ${\sf k}$ is a field. By [48, Theorem 1.9.(a)(ii)] the module $Q(e_{w})$ is ${\mathrm{ind}}_{\mathbf{B}}^{\mathbf{P}}$ acyclic. Lemma 13.9, together with the Generalized Tensor Identity [32, I Proposition 4.8], give
$\mathrm{R}^{j}{\mathrm{ind}}_{\mathbf{B}}^{\mathbf{P}}(P(-e_{v})\otimes Q(e_{w}))=P(-e_{v})\otimes\mathrm{R}^{j}{\mathrm{ind}}_{\mathbf{B}}^{\mathbf{P}}(Q(e_{w}))=\begin{cases}P(-e_{v})\otimes\hat{Q}(e_{w})&\text{if $j=0$,}\\ 0&\text{else.}\end{cases}$
So $\mathrm{R}^{i}{\mathrm{ind}}_{\mathbf{B}}^{\mathbf{G}}(P(-e_{v})\otimes Q(e_{w}))=\mathrm{R}^{i}{\mathrm{ind}}_{\mathbf{P}}^{\mathbf{G}}(P(-e_{v})\otimes\hat{Q}(e_{w}))$. But the left hand side is known from Theorem 6.6. ∎
Proof of Theorems 13.2, 13.3, 13.5.
With these replacements in hand, we now proceed as in the case $\mathbf{P}=\mathbf{B}$: For $p\in W^{\mathbf{P}}$, let $\hat{X}_{p}$ be the image of $P(-e_{p})^{*}$ under the left adjoint of the inclusion of $\hat{\sf Q}_{\succeq p}$ into $\mathcal{D}=\mathop{\mathrm{D}\kern 0.0pt}\nolimits^{b}({\mathrm{rep}}(\mathbf{P}))$. And let $\hat{Y}_{p}$ be the image of $\hat{Q}(e_{p})$ under the right adjoint of the inclusion of $\hat{\sf P}_{\preceq p}$ into $\mathcal{D}$.
Let $f_{p}$ be the natural map from $P(-e_{p})^{*}$ to $Q(e_{p})$. Recall that $f_{p}$ factors as a surjection $P(-e_{p})^{*}\to{\sf k}_{e_{p}}$, followed by an injection ${\sf k}_{e_{p}}\to Q(e_{p})$. Let $\hat{f}_{p}:P(-e_{p})^{*}\to\hat{Q}(e_{p})$ be induced by $f_{p}$. Then $\hat{f}_{p}$ induces an isomorphism of $\mathbf{T}$-modules $(P(-e_{p})^{*})_{e_{p}}\to(\hat{Q}(e_{p}))_{e_{p}}$ and the weights $\mu$ of ${\mathsf{ker}}(\hat{f}_{p})$, ${\mathsf{coker}}(\hat{f}_{p})$ satisfy $\mu<_{a}e_{p}$. We have an exact triangle